be give 1. cos(135°) 2. tan 3.sin(-5) 4. sec(-210°)

no cal
Answer #1

And we have to bind cos (135°) Cos (135)= cos(185-45) AS COS(190-0)=- woso = cos(135) = ws 45° Cos(135) 2 Hence ws (135) = AAns sec (-210) As sec (-0) =+seco 5 sec (-2100) = sec (2100) - 0 sec (210) = sec (1909+30°) As sec (180°+0) = - seco see (210